OpenStudy (anonymous):

can you use synthetic division to show that x = -4 is a root of this equation

OpenStudy (anonymous):

use synthetic division to show that x = -4 is a root of the equation x^3 + 1x^2 - 30x - 72 = 0

OpenStudy (anonymous):

Then, use the result to factor the polynomial completely into the form (x+A)(x+B)(x+C) Then give the list of values A,B,C you obtain:

OpenStudy (anonymous):

you can check that -4 is a zero by substitution. then factor as \[(x+4)\times (\text{something)}\] and you can get the something by division

OpenStudy (anonymous):

it will be \[(x+4) (x-6) (x+3) = 0\]